OpenStudy (anonymous):

Let T: P1---> R^2 be the function defined by the formula T(P0,P2) a) find T (2-x) b) find T^-1 Please, help

OpenStudy (anonymous):

problem 17. a little bit different when in review set, it turns T(P0,P2)

OpenStudy (anonymous):

\[T(1-2x)=(1,-1)\]

OpenStudy (anonymous):

T x means?

OpenStudy (anonymous):

\[p(x)=1-2x\]

OpenStudy (anonymous):

p(x) = a0 + a1 *x

OpenStudy (anonymous):

yup and you are given \(a_0,a_1\)

OpenStudy (anonymous):

look at your typo!!! it is so nonsense!! P (0) = a0 P(1) = a0 + a1 P(2) = a0 + a1*x

OpenStudy (anonymous):

prob 17, right?

OpenStudy (anonymous):

yes, but I am working with the same problem a little bit different from #17. That problem define T (Px)= (P0,P1) , Mine T (P(x)) = (P0,P2)

OpenStudy (anonymous):

ok.. \[T(p(x))=(p(0),p(1))\\ {\rm when}\;p(x)=2-x\\ T(2-x)=\left(2-(0),2-(1)\right)=\mathbf{(2,1)} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good, so, what is the question now?

OpenStudy (anonymous):

find T^-1

OpenStudy (anonymous):

ok, sorry, I got it.

OpenStudy (anonymous):

\[ {\rm let}\;T^{-1}(x)=a_0+a_1x\\ p(0)=a_0\qquad p(1)=p(0)+a_1\implies a_1=p(1)-p(0) \]

OpenStudy (anonymous):

kid, a chain of letter, how can I know what is what?

OpenStudy (anonymous):

I did not understadn that.

OpenStudy (anonymous):

I don't know what's wrong with my computer, let me copy what I see in your post T x a a x p a p p a a p p to me, it is no meaning

OpenStudy (anonymous):

hold on. I think I am wrong with the inverse thing

OpenStudy (anonymous):

It should also be another matrix.

OpenStudy (anonymous):

ooh.. try refreshing the page

OpenStudy (anonymous):

press the "F5" button on the keyboard

OpenStudy (anonymous):

Ok, it makes sense now

OpenStudy (anonymous):

I got it, kid.

OpenStudy (anonymous):

the inverse too?

OpenStudy (anonymous):

no, just got how to get T =(2,1)

OpenStudy (anonymous):

ok for inverse, we'd need the p(x)

OpenStudy (anonymous):

and we find the standard matrix for the transformation

OpenStudy (anonymous):

@Hoa you there?

OpenStudy (anonymous):

yes, sir

OpenStudy (anonymous):

so, are we taking \(p(x)=2-x\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then\[T(p(x))=(2,1)\] and its standard matrix is non-invertible!!

OpenStudy (anonymous):

since the linear mapping is between two regions of different dimensions

OpenStudy (anonymous):

is it not T^-1(2,1) = P(x)?

OpenStudy (anonymous):

thats the only way I can think of it.

OpenStudy (anonymous):



OpenStudy (anonymous):

but, in general, you have an inverse for \(\mathbb{R}^n\rightarrow\mathbb{R}^n\)

OpenStudy (anonymous):

I know what you mean. ok, so the conclusion is T is non-invertible matrix?

OpenStudy (anonymous):

not "T" but the inverse matrix of the transformation does not exist. T -> is the transformation, not the matrix -> has a system matrix "A" whose inverse may or maynot exist. T^{-1} -> is the inverse transformation